Write the slope-intercept equation for the line that passes through (-4, -3) and (-6, 7).
Answer by John10(297)

Write the slope-intercept equation for the line that passes through (-4, -3) and (-6, 7).
when two points are given, we use slope form to find slope
m = (y2- y1)/(x2 - x1) = (7 + 3)/(-6+4) = 10/ -2 = -5
Then use the found slope and one of the given points
y - yo = m(x -xo) where (xo,yo) is the given point
y - 7 = -5 (x + 6)
y - 7 =-5 -30
y = -5x - 23
The equation is already in slope-intercept form!!!!!!!!